OpenStudy (anonymous):

If f(x) = x + 4, find the value of x if f(x) = 6. A. 6 B. 4 C. 2 D. 10

OpenStudy (anonymous):

Set x + 4 equal to six and solve for x.

OpenStudy (anonymous):

Is it 10?

OpenStudy (anonymous):

No. x + 4 = 6. to get x by itself on the left, you must subtract 4 from each side.

OpenStudy (anonymous):

So, is it 6?

OpenStudy (anonymous):

What's six minus four?

OpenStudy (anonymous):

f(6) = x + 4 = 6 x + 4 -4 = 6 - 4 x + 0 = 2 x = 2

OpenStudy (anonymous):

Yeah, I got it like five minutes ago. But I didn't say anything 'cause you weren't here at the time... Thanks.